# Black-Hole Cartography

### v1

@@ Appendix B Simultaneous mapping of multiple QNMs
Appendix C Numerical error
The numerical error, which is expected to be present in all NR simulations, is determined in several ways throughout this paper. All the methods involve comparing the preferred version of the CCE simulations (Level 5, second smallest radius) to the other possible combinations of the Level 4 and 5, and second and third smallest worldtube radii.
For the plots in Fig. 2, the three other possible waveforms are each shifted in time to find the optimal alignment that minimizes the mismatch with the preferred waveform. This is done by rolling the waveform by discrete time intervals and computing the mismatch until the minimum is found, then shifting the position of $t=0$ by this corresponding amount. After this alignment, the mismatch was calculated between the preferred simulation and the three remaining simulations from $t_{0}$ to $T=t_{0}+100M$. The maximum mismatch was taken at each $t_{0}$ and used as the numerical error at this point. This is the approach given in [25].
For the spatial mismatch plots in Fig. 3, the mapping was performed with the four simulation level and radius combinations. The spatial mismatch was calculated between the reconstruction from the preferred simulation and each of the remaining three and the maximum value at each $t_{0}$ was used to quantify the numerical error.
In Figs 4, 5, 9, and 10 where a particular numerical value is recovered, a range was computed by determining the value for all simulation level and radius combinations and plotting the maximum and minimum at each $t_{0}$ as a translucent region around the preferred simulation value.
@@ Appendix C Numerical error
Figure 14: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the QNM $(2,2,0,+)$ are compared for two no-spin, equal mass ratio progenitor simulations. The bold lines are obtained using the CCE simulation 0001, and the dotted lines are obtained from simulation 0001 in the catalog of extrapolated waveforms.
Appendix D Computing quadratic structure predictions
@@
where the mode-mixing coefficients are defined by the following integral,
@@
This is consistent with the definition in Eq. 4, but with the spin-weight dependence of the mode-mixing coefficients made explicit in the notation. Note that the mode-mixing coefficients are not spin weight $s$ objects; here, the preceding subscript $s$ is intended as a reminder that these coefficients are to be used in the expansion of the spin-weight $s$ spheroidal harmonics. The mode-mixing coefficients are functions of the BH remnant spin $\chi_{f}$ (although this dependence is suppressed in our notation for clarity) and are computed using the qnm package.
In a similar way, the coefficients ${}_{ss^{\prime}}\kappa^{\beta\beta^{\prime}\beta^{\prime\prime}}$ allow us to re-expand products of two spin-weighted spherical harmonics in the spin-weighted spherical harmonics basis. Because the product of a spin-weight $s$ and a spin weight $s^{\prime}$ field has spin weight $s+s^{\prime}$, this re-expansion is done in terms of the spherical harmonics with this spin weight,
@@
Similar expressions can be written down for the other spin-weight combinations: B (ii) and so on.
Prediction C. It is difficult to write down this prediction concisely in the same notation used in this section. Instead, it is more convenient to compute the mixing directly from Eq. 30 using spheroidal harmonics with modified spheroidicity determined using the spheroidal package.
Prediction D. The spin-weight raising operator satisfies ð$\eth{}_{s}Y^{\beta}(\theta,\phi)=\sqrt{(\beta_{\ell}-s)(\beta_{\ell}+s+1)}{}_{s+1}Y^{\beta}(\theta,\phi)$, where $\beta_{\ell}$ is the polar $\ell$ index from the $\beta$ tuple of spherical indices. From Eq. 22, the term inside the parentheses is identical to prediction A and can be written as Eq. 33, then acting twice with the spin-weight raising operator gives,

### v2

@@ Appendix C Simultaneous mapping of multiple QNMs
Appendix D Numerical error
The numerical error, which is expected to be present in all NR simulations, is determined in several ways throughout this paper. All the methods involve comparing the preferred version of the CCE simulations (Level 5, second smallest radius) to the other possible combinations of the Levels 4 and 5, and the second and third smallest worldtube radii.
For the plots in Fig. 2, the three other possible waveforms are each shifted in time to find the optimal alignment that minimizes the mismatch with the preferred waveform. This is done by rolling the waveform by discrete time intervals and computing the mismatch until the minimum is found, then shifting the position of $t=0$ by this corresponding amount. After this alignment, the mismatch is calculated between the preferred simulation and the three remaining simulations from $t_{0}$ to $T=t_{0}+100M$. The maximum mismatch is taken at each $t_{0}$ and used as the numerical error at this point. This is the approach given in [25].
For the spatial mismatch plots in Fig. 3, the mapping is performed with the four simulation level and radius combinations. The spatial mismatch is calculated between the reconstruction from the preferred simulation and each of the remaining three, and the maximum value at each $t_{0}$ is used to quantify the numerical error.
In Figs 4, 5, 9, and 10, where a particular numerical value is recovered, a range is computed by determining the value for all simulation level and radius combinations and plotting the maximum and minimum at each $t_{0}$ as a translucent region around the preferred simulation value.
@@ Appendix D Numerical error
Figure 15: The magnitude of the decay-corrected amplitudes obtained from spatial mapping of the QNM $(2,2,0,+)$ are compared for two no-spin, equal-mass-ratio progenitor simulations. The bold lines are obtained using the CCE simulation 0001, and the dotted lines are obtained from simulation 0001 in the catalog of extrapolated waveforms.
Appendix E Computing quadratic structure predictions
@@
where the mode-mixing coefficients are defined by the following integral:
@@
This is consistent with the definition in Eq. 4, but with the spin-weight dependence of the mode-mixing coefficients made explicit in the notation. Note that the mode-mixing coefficients are not spin-weight $s$ objects; here, the preceding subscript $s$ is intended as a reminder that these coefficients are to be used in the expansion of the spin-weight $s$ spheroidal harmonics. The mode-mixing coefficients are functions of the BH remnant spin $\chi_{f}$ (although this dependence is suppressed in our notation for clarity) and are computed using the qnm package.
In a similar way, the coefficients ${}_{ss^{\prime}}\kappa^{\beta\beta^{\prime}\beta^{\prime\prime}}$ allow us to re-expand products of two spin-weighted spherical harmonics in the spin-weighted spherical harmonics basis. Because the product of a spin-weight $s$ and a spin weight $s^{\prime}$ field has spin-weight $s+s^{\prime}$, this re-expansion is done in terms of the spherical harmonics with this spin weight,
@@
Similar expressions can be written for the other spin-weight combinations – B (ii) and so on.
Prediction C. It is difficult to write this prediction concisely in the same notation used in this section. Instead, it is more convenient to compute the mixing directly from Eq. 30 using spheroidal harmonics with modified spheroidicity, determined using the spheroidal package.
Prediction D. The spin-weight raising operator satisfies ð$\eth{}_{s}Y^{\beta}(\theta,\phi)=\sqrt{(\beta_{\ell}-s)(\beta_{\ell}+s+1)}{}_{s+1}Y^{\beta}(\theta,\phi)$, where $\beta_{\ell}$ is the polar $\ell$ index from the $\beta$ tuple of spherical indices. From Eq. 22, the term inside the parentheses is identical to prediction A and can be written as Eq. 33, then acting twice with the spin-weight raising operator gives